OpenStudy (anonymous):

Solve the following: LOG (8x+1) = 1 + LOG (X-4) I know the answer is 41/2. Just looking for the steps.

myininaya (myininaya):

log(8x+1)-log(x-4)=1 log([8x+1]/[x-4])=1 10^ 10^ (8x+1)/(x-4)=10 8x+1=10(x-4) 8x+1=10x-40 1+40=10x-8x 41=2x x=41/2

OpenStudy (anonymous):

Alternate solution: \[\log (8x+1)=1+\log(x-4) \] \[\log (8x+1)=\log(10) +\log(x-4) \] \[\log (8x+1)=\log(10(x-4)) \] \[8x+1=10x-40\] \[41=2x\] \[x={41 \over 2}\]